# HidePrint: Hiding the Radio Fingerprint via Random Noise

### v1

@@ 7.1 Wired Link - Adversary $\mathcal{A}_{1}$
Training on noise-free samples. We first consider the measurements acquired using the wired communication link and focus on training the ResNet-18 CNN with noise-free data and testing on any noise level. As previously introduced in Sect. 5, for each device, we collected five different noise levels, i.e., standard deviation $\sigma=\{0,0.01,0.02,0.03,0.04,0.05\}$, considering five different noise models, i.e., noise-free, Gaussian, Impulse, Laplacian, and Uniform. Therefore, we considered 10 classes (one for each transmitter) and trained a model using the 10 available measurements with noise levels equal to zero. Subsequently, we tested our model considering all combinations of transmitter and noise levels (greater than zero). Figure 5 shows the results of our analysis. In particular, the upper part of Fig. 5 depicts the accuracy of the classifier as a function of the noise level ($\sigma$) while considering all the noise types provided by the GnuRadio block Noise Source. We observe that the classification accuracy is not affected by the noise type, since they all exhibit the same behavior. In contrast, the accuracy is strongly affected by the noise variance $\sigma$, and that $\sigma\approx 0.02$ is enough to achieve the random guess (accuracy equal to 0.1). In the bottom part of Fig. 5, we considered the SNR as a function of the noise level introduced by the transmitter. We would like to highlight that a variation in the SNR of about 0.1dB is enough to achieve anonymity, thus proving that the quality of the link is negligibly affected by the injection of noise. For the sake of completeness, we recall Fig. 2, which has been generated by considering $10^{5}$ samples from the noise-free configuration and $10^{5}$ samples from the configuration combining Gaussian noise and $\sigma=0.05$.
@@
Figure 5: Adversary $\mathcal{A}_{1}$ and training on no-noise samples. The adversary trains a model with 10 classes (one for each transmitter) on measurements with no noise and then is challenged to identify the transmitter by considering measurements with noise strictly greater than zero.
@@
Finally, we consider the t-Distributed Stochastic Neighbor Embedding (t-SNE) analysis associated with two cases, i.e., noise-free and $\sigma=0.02$, as shown in Fig. 6. The t-SNE analysis performs a dimensionality reduction, i.e., from 10 activations extracted from the last layer of the ResNet-18 network to a two-dimensional space that can be represented in a Cartesian plane. As can be seen from Fig. 6(a), noise-free values are very well disjoint, and clustering makes it possible to assign each point (image) to the related class. This is not possible in the scenario of Fig. 6(b), where $\sigma=0.02$ leads the images coming from the 10 different classes to overlap.
@@
Figure 6: t-SNE analysis for two reference cases: noise-free and $\sigma=0.02$. Classification is possible for the noise-free scenario, but adding noise makes the transmitters virtually indistinguishable.
Training on all samples. We also consider the scenario where the adversary trains the ResNet-18 CNN on all the (10) transmitters in the pool and noise levels while considering only the Gaussian noise model—recall from Fig. 5 that different types of noise have the same impact on the fingerprint. To achieve this, we consider the whole dataset and split it into three disjoint chunks, i.e., 60%, 20%, and 20%, for training, validation, and testing, respectively. This adversary configuration is the most powerful considered in this work. Indeed, as the most favorable condition for the adversary, we assume that the adversary is able to collect (and label) samples from all the transmitters with all the noise levels. Figure 7(a) shows the confusion matrix associated with the actual and predicted transmitters. We observe that, with such knowledge, $\mathcal{A}_{1}$ significantly increases its chances of detecting the presence of a transmitter of the pool in the radio spectrum (the average accuracy is about 0.66). Nevertheless, our solution still hides the identity of the transmitters, as depicted in Figure 7(b). Indeed, the average False Positive Ratio (FPR) and False Negative Ratio (FNR), computed on all the transmitters, are similar and equal to 0.35 (dashed red and blue line in Fig. 7(b)). A significant number of transmitters (6 out of 10) are above the average, meaning that when the classifier predicts a specific transmitter, the probability of an error is higher than 0.35. We stress that this value has been obtained by giving to the adversary all the possible knowledge available in the system.
@@
Figure 7: Adversary $\mathcal{A}_{1}$ and training on all noise levels. The adversary trains a model with 10 classes (one for each transmitter) on measurements with all the noise levels spanning between 0 and 0.05, and then it is challenged to identify the transmitter by considering random samples taken from random measurements with noise strictly greater than zero.
7.2 Wired Link - Adversary $\mathcal{A}_{2}$
@@
Training on noise-free samples. We still consider the measurements acquired using the wired link and consider the case of training on noise-free samples for a specific device (only), in line with the adversary model $\mathcal{A}_{2}$. Then, we tested on all the available transmitter configurations, i.e., noise-free and with Gaussian noise with standard deviation $\sigma=\{0,0.01,0.02,0.03,0.04,0.05\}$. We highlight that we did consider only one noise model (Gaussian) in this case since the noise pattern does not affect the accuracy of the classification process (recall Sec. 7.1 and Fig. 5). Figure 8 shows the results of our analysis in terms of MSE as a function of the transmitter identifier (ID) used for training. Each bar represents the quantiles 5, 50, and 95 associated with the MSE computed on the images generated from the IQ samples. We highlight the comparisons we have made considering four colors: (i) the comparison of the reference set, i.e., two disjoint datasets (60% and 40%) taken from each transmitter ID (red color), one used for training and another one used for testing, respectively; (ii) the comparison of the reference transmitter ID and any other transmitter ID in the pool while considering noise-free samples (blue color), (iii) the comparison of each transmitter ID with the same ID but considering the measurements with the Gaussian noise (green color), and finally, (iv) the comparison of the reference transmitter ID with any other transmitters but considering the measurements with the noise (black color).
@@
Figure 8: Adversary $\mathcal{A}_{2}$ and training on noise-free samples. The adversary trains 10 different models, one for each transmitter considering only the noise-free samples. Red bars represent the test results performed on the same measurements considered for training (considering a split ratio of 80/20). Subsequently, the adversary challenges the trained model with the remainder of the measurement in the dataset: (i) noise-free samples from different transmitters (blue bars), (ii) noisy samples from the same transmitter considered for the training (green bars), and finally, (iii) noisy samples from different transmitters (black bars).
@@
The red bars in Fig. 8 serve as the reference values. We selected 10 noise-free measurements from each transmitter and then split the generated images into two datasets (60% and 40%) for training and testing, respectively. This setup provides the MSE reference values to be compared with the images coming from other configurations. When comparing the red and blue bars (different transmitter IDs and noise-free measurements), we do not see any (major) overlap that proves that transmitters can be distinguished in a noise-free environment. The minor overlap between the transmitters 2 and 9 will be investigated later on. An orthogonal analysis involves the performance estimation of our solution by comparing the noise-free measurements (red bars) with the ones affected by HidePrint (green bars), where we injected Gaussian noise with different standard deviations. Note that the noise significantly affects the MSE, and in particular, it makes blue bars overlap with the green ones, i.e., the reference transmitter becomes indistinguishable from the others when adding noise to the transmitted signal. Finally, the black error bars consider the measurements with noise from all the transmitters. Note that the noise makes all the measurement configurations overlap, thus achieving indistinguishability between the reference transmitter and the others in the pool.
@@
We now quantify the accuracy that can be achieved considering the attack deployed by adversary $\mathcal{A}_{2}$. For each transmitter, we define a decision threshold according to Eq. 6 defined earlier. Figure 9 shows the results of our analysis while comparing noise-free signals (green shaded areas) with the remainder of our dataset (red shaded areas). We highlight in blue the thresholds computed for each transmitter. Our analysis shows that the fingerprint (completely) changes when noise is added to the transmitted signal. Indeed, the false positive ratio, i.e., samples from noise-free measurements leading to MSE values larger than the threshold values, is zero, independently of the considered transmitter. Moreover, the false negative ratio, i.e., samples taken from any other measures, is always zero except for two cases, which are about 0.01. This is the case of the transmitters 2 and 9, which have a similar fingerprint (recall Fig. 8). Finally, recalling Fig. 8, we observe that different configurations (noise-free and noisy) eventually collapse in the same range (red shaded areas in Fig. 9), thus making the detection of each of them challenging. For example, an adversary might be able to collect noisy samples with a specific noise level, then train a model, and finally accept the challenge of detecting such a pattern in the radio spectrum. This is unlikely, since all comparisons fall in the region with MSE values between 11 and 19, and therefore, the adversary would experience a high number of misclassifications independently of the chosen configuration.
@@
Figure 9: Comparison between MSE values when training on noise-free samples (green curves) and testing on the remainder of the dataset (red curves). We highlight with a blue line the thresholds computed according to Eq. 6. The overlap between the two sets of distributions (red and blue) is negligible for all the transmitters. The autoencoder is successful in distinguishing the transmitters.
@@
Training on all samples. In the following, we consider the case of adversary $\mathcal{A}_{2}$ owning a model trained on both noise-free and noisy samples. Indeed, the adversary has been exposed to the full dataset (noise-free and noisy samples) associated with each device considered in our test. As for the previous case, we considered the Gaussian noise with standard deviation $\sigma=\{0,0.01,0.02,0.03,0.04,0.05\}$. Figure 10 shows the results of our analysis, where we perform a targeted attack (adversary $\mathcal{A}_{2}$) on each device of the pool (transmitter IDs). We report, for each of them, the MSE associated with the test set from the selected transmitter ID (green shaded area) and the MSE associated with the other measurements (other transmitters with any noise level). The probability distribution functions overlap for all the transmitter IDs. When considering the threshold defined by Eq. 6 (blue lines in Fig. 10), we obtain an average false negative ratio of about 0.66, confirming that the autoencoder fails to detect the presence of a specific transmitter in the radio spectrum when the training is performed on all the available samples (noise-free and noisy). Such a finding is confirmed by the empirical intuition that a model trained on all the samples cannot differentiate a specific transmitter from the pool since the (injected) noise makes the fingerprint of all the transmitters very similar to each other.
@@
Figure 10: MSE values when training on any samples (green curves) and testing on the remainder of the dataset (red curves). We highlight with a blue line the thresholds computed according to Eq. 6. The overlap between the two sets of distributions (red and blue) is not negligible for all the transmitters (0.66 on average), and thus, the autoencoder fails to distinguish the transmitters.
7.3 Wireless Link
@@
Our wireless scenario involves the same setup and configuration as the wired one (recall Fig. 3). The transmitters and the receiver have been placed a few meters far away in an office during working time, i.e., with people moving around in close proximity to the transmitter-receiver link, generating multipath. We consider the same data collection procedure as that described for Fig. 5, i.e., Gaussian noise type, CNN ResNet-18, and $10^{5}$ samples per image. Figure 11 (top) shows the accuracy of ResNet-18 as a function of the noise standard deviation $\sigma$. When $\sigma=0$ (noise-free), the accuracy is approximately 0.96, while dropping to 0.16 when $\sigma=0.02$. It is worth noting that an accuracy value of 0.1 is equivalent to a random guess. The SNR is consistent with the one experienced when using the cable: we kept the attenuation to 30dB and the distance between the transmitter and the receiver was only a few meters. Moreover, we observe the existence of a transient region in the range ]0,0.02]]0,0.02]] 0 , 0.02 ] where the accuracy is affected by the combination of different effects, i.e., the fingerprint of the device and the fingerprint of the noise introduced by the channel [27, 26]. The measurements involving the RF cable are not affected by external factors (multipath), and therefore, the fingerprint of the device is the principal component of the received signal. In contrast, when transmitting over the wireless channel, the channel impulse response varies over time depending on the communication link. Therefore, the DL model is likely to experience either consistent or different channel states (channel fingerprint). The final result, as observed by other contributions [27, 26, 8], is that RFF is affected by both the device and the channel fingerprint. The accuracy for $\sigma=0.01$ in Fig. 11 (top), being equal to 0.56, is higher than the one in Fig. 5 (top), being equal to 0.25, and it can be explained considering the impact of the channel fingerprint in the overall RFF process: the fingerprint of the wireless channel has a higher impact than the one in the cable. Indeed, the measurements taken considering different transmitters have the same cable with the same channel fingerprint, while this is not happening in the radio scenario, where different transmitters might be affected by a (slightly) different radio channel, which helps the classification. Finally, we observe that a noise standard deviation of 0.02 (and higher) can fully anonymize the transmitter independently of the considered measurement scenario. Moreover, the drop associated with the SNR (0.1dB) necessary to remove the fingerprint is consistent with the one experienced using the cable.
@@
Figure 11: Adversary $\mathcal{A}_{1}$ training on noise-free samples from a wireless link. The adversary trains a model with 10 classes (one for each transmitter) on measurements with noise-free samples and then has to identify the transmitter(s) by considering measurements with noise strictly greater than zero.
@@
We now consider adversary $\mathcal{A}_{2}$, i.e., a targeted attack against a specific transmitter in our pool of 10. We trained an autoencoder on wireless measurements taken from a specific device while setting the noise level $\sigma$ and tested it on the remainder of the measurements collected from the other transmitters in the pool given the same $\sigma$. In fact, we assume $\mathcal{A}_{2}$ is capable of training from any type of measurement (noise-free and noisy) and then test it in the wild by collecting samples from the radio spectrum. Figure 12 shows the results of our analysis, where we consider the average Receiver Operating Characteristic (ROC) curves calculated considering the mean value over the 10 transmitters when varying $\sigma\in\{0,0.01,0.02,0.03,0.04,0.05\}$. Our results confirm our previous findings. First, we observe that the performance of the autoencoder in a noise-free environment (solid blue line in Fig. 12)) is characterized by an FPR and TPR of 0.21 and 0.78, respectively, when considering the point in the ROC curve that offers the best trade-off between sensitivity and specificity, under the assumption that both types of classification errors (TPR and FPR) cost the same—the overall accuracy is equal to 0.81. Moreover, a Gaussian noise with standard deviation $\sigma=0.01$ is not enough to de-anonymize the transmitter. Indeed, the solid red line in Fig. 12, representing the ROC curve for $\sigma=0.01$, is characterized by FPR and TPR equal to about 0.27 and 0.7, respectively—the overall accuracy is equal to 0.73. Similarly to the previous case ($\mathcal{A}_{1}$), $\sigma=0.02$ allows us to achieve a much better anonymization level, i.e., TPR and FPR values equal to 0.27 and 0.7, respectively, with an overall accuracy equal to 0.64.
@@
Figure 12: Adversary $\mathcal{A}_{2}$. We trained an autoencoder on measurements taken from a target device and tested against the others in the pool while keeping the same noise level $\sigma\in\{0,0.01,0.02,0.03,0.04,0.05\}$ in an actual wireless link.
8 Comparison with other RFF solutions

### v2

@@ 7.1. Wired Link - Adversary $\mathcal{A}_{1}$
Training on noise-free samples. We first consider the measurements acquired using the wired communication link and focus on training the ResNet-18 CNN with noise-free data and testing on any noise level. As previously introduced in Sect. 5, for each device, we collected six different noise levels, i.e., standard deviation $\sigma=\{0,0.01,0.02,0.03,0.04,0.05\}$, considering five different noise models, i.e., noise-free, Gaussian, Impulse, Laplacian, and Uniform. We observe that $\sigma$ is dimensionless, since both I and Q components are normalized to be placed at the ideal positions in the IQ plane as per Fig. 2. Therefore, we considered 10 classes (one for each transmitter) and trained a model using the 10 available measurements with noise levels equal to zero. Subsequently, we tested our model considering all combinations of transmitters and noise levels (greater than zero). Figure 5 shows the results of our analysis. In particular, the upper part of Fig. 5 depicts the accuracy of the classifier as a function of the noise level ($\sigma$) while considering all the noise types provided by the GNURadio block Noise Source. We observe that the classification accuracy is not affected by the noise type, since they all exhibit the same behavior. In contrast, the accuracy is strongly affected by the noise standard deviation $\sigma$. Starting from accuracy $0.99$ with $\sigma=0$ (baseline), the value $\sigma\approx 0.02$ is enough to achieve the random guess (accuracy close to 0.1). In the bottom part of Fig. 5, we considered the SNR as a function of the noise level introduced by the transmitter. We highlight that a variation in the SNR of about 0.1dB is enough to achieve anonymity, thus proving that the quality of the link is negligibly affected by the injection of noise. For the sake of completeness, we recall Fig. 2, which has been generated by considering $10^{5}$ samples from the noise-free configuration and $10^{5}$ samples from the configuration combining Gaussian noise and $\sigma=0.05$.
@@
Figure 5. Adversary $\mathcal{A}_{1}$ training on noise-free samples and testing on measurements with noise greater than zero. The adversary trains a model with 10 classes (one for each transmitter) on measurements with no noise and then is challenged to identify the transmitter.
@@
Finally, we consider the t-Distributed Stochastic Neighbor Embedding (t-SNE) analysis associated with two cases, i.e., noise-free and $\sigma=0.02$, as shown in Fig. 6. The t-SNE analysis performs a dimensionality reduction, i.e., from 10 activations extracted from the last layer of the ResNet-18 network to a two-dimensional space that can be represented in a Cartesian plane. As shown in Fig. 6(a), noise-free values are very well disjoint, and clustering makes it possible to assign each point (image) to the related class. This is not possible in the scenario of Fig. 6(b), where $\sigma=0.02$ leads to the overlap of the images of the different classes.
@@
Figure 6. t-SNE analysis for noise-free and $\sigma=0.02$. Classification is possible with no noise, but adding noise makes the transmitters virtually indistinguishable when $\sigma\geq 0.02$.
Training on all samples. We also consider the scenario where the adversary trains the ResNet-18 CNN on all the (10) transmitters in the pool and noise levels while considering only the Gaussian noise model—recall from Fig. 5 that different types of noise have the same impact on the fingerprint. To achieve this, we consider the whole dataset and split it into three disjoint chunks, i.e., 60%, 20%, and 20%, for training, validation, and testing, respectively. This adversary configuration is the most powerful considered in this work. Indeed, as the most favorable condition for the adversary, we assume that the adversary is able to collect (and label) samples from all the transmitters with all the noise levels. Figure 7(a) shows the confusion matrix associated with the actual and predicted transmitters. We observe that, with such knowledge, $\mathcal{A}_{1}$ significantly increases its chances of detecting the presence of a transmitter of the pool in the radio spectrum (the average accuracy is about 0.66). Nevertheless, our solution still hides the identity of the transmitters, as depicted in Figure 7(b). Indeed, the average False Positive Ratio (FPR) and False Negative Ratio (FNR), computed on all the transmitters, are similar and equal to 0.35 (dashed red and blue line in Fig. 7(b)). A significant number of transmitters (6 out of 10) are above the average, meaning that when the classifier predicts a specific transmitter, the probability of an error is higher than 0.35. We stress that this value has been obtained by giving to the adversary all the possible knowledge available in the system. Moreover, recalling that the transmitter and the receiver are connected by an RF cable, we highlight that this setup avoids the multipath fading, which in turn, helps as well to hide the fingerprint, thus representing an upper bound on the performance of the adversary.
@@
Figure 7. Adversary $\mathcal{A}_{1}$ training on all noise levels and testing with random samples taken from random measurements with noise strictly greater than zero. The adversary trains a model with 10 classes (one for each transmitter) on measurements with all the noise levels spanning between 0 and 0.05, and then it is challenged to identify the transmitter.
7.2. Wired Link - Adversary $\mathcal{A}_{2}$
@@
Training on noise-free samples. We still consider the measurements acquired using the wired link and consider the case of training on noise-free samples for a specific device (only), in line with the adversary model $\mathcal{A}_{2}$. Then, we tested on all the available transmitter configurations, i.e., noise-free and with Gaussian noise with standard deviation $\sigma=\{0,0.01,0.02,0.03,0.04,0.05\}$. We highlight that we did consider only one noise model (Gaussian) in this case since the noise pattern does not affect the accuracy of the classification process (recall Sec. 7.1 and Fig. 5). Figure 8 shows the results of our analysis in terms of MSE as a function of the transmitter identifier (ID) used for training. Each bar represents the quantiles 5, 50, and 95 associated with the MSE computed on the images generated from the IQ samples. We highlight the comparisons we have made considering four colors: (i) the comparison of the reference set, i.e., two disjoint datasets (60% and 40%) taken from each transmitter ID (red color), one used for training and another one used for testing, respectively; (ii) the comparison of the reference transmitter ID and any other transmitter ID in the pool while considering noise-free samples (blue color), (iii) the comparison of each transmitter ID with the same ID but considering the measurements with the Gaussian noise (green color), and finally, (iv) the comparison of the reference transmitter ID with any other transmitters but considering the measurements with the noise (black color). The red bars in Fig. 8 serve as the reference values. We selected 10 noise-free measurements from each transmitter and then split the generated images into two datasets (60% and 40%) for training and testing, respectively. This setup provides the MSE reference values to be compared with the images coming from other configurations.
@@
Figure 8. The adversary $\mathcal{A}_{2}$ trains 10 different models, one for each transmitter, considering only the noise-free samples. Red bars represent the test results performed on the same measurements considered for training (split ratio 80/20). Subsequently, the adversary challenges the trained model with additional datasets: (i) noise-free samples from different transmitters (blue bars), (ii) noisy samples from the same transmitter considered for the training (green bars), and finally, (iii) noisy samples from different transmitters (black bars).
@@
When comparing the red and blue bars (different transmitter IDs and noise-free measurements), there is no (major) overlap that proves that transmitters can be distinguished in a noise-free environment. The minor overlap between transmitters 2 and 9 will be investigated later on. An orthogonal analysis involves the performance estimation of our solution by comparing the noise-free measurements (red bars) with the ones affected by HidePrint (green bars), where we injected Gaussian noise with different standard deviations. The injected noise significantly affects the MSE, and in particular, it makes blue bars overlap with the green ones, i.e., the reference transmitter becomes indistinguishable from the others. Finally, the black error bars represent the measurements with noise from all the transmitters. Note that the noise makes all the measurement configurations overlap, thus achieving transmitter indistinguishability.
@@
To quantify the accuracy that can be achieved considering the attack deployed by adversary $\mathcal{A}_{2}$, for each transmitter, we define a decision threshold according to Eq. 6. Figure 9 shows the results of our analysis while comparing noise-free signals (green shaded areas) with the remainder of our dataset (red shaded areas). We highlight in blue the thresholds computed for each transmitter. Our analysis shows that the fingerprint (completely) changes when noise is added to the transmitted signal. Indeed, the FPR, i.e., samples from noise-free measurements leading to MSE values larger than the threshold values, is zero independently of the considered transmitter. The false negative ratio, i.e., samples taken from any other measures, is always zero except for two cases, which are about 0.01. This is the case of the transmitters 2 and 9, which have a similar fingerprint (recall Fig. 8). Finally, recalling Fig. 8, we observe that different configurations (noise-free and noisy samples) eventually collapse in the same range (red shaded areas in Fig. 9), thus making the detection of each of them challenging. For example, an adversary might be able to collect noisy samples with a specific noise level, then train a model, and finally accept the challenge of detecting such a pattern in the radio spectrum. This is unlikely, since all comparisons fall in the region with MSE values between 11 and 19, and therefore, the adversary would experience a high number of misclassifications independently of the chosen configuration.
@@
Figure 9. MSE values when training on noise-free samples (green curves) and testing on the rest of the dataset (red curves). The blue lines are the thresholds computed via Eq. 6. The overlap between the two sets of distributions (red and blue) is negligible for all the transmitters: the autoencoder is successful in distinguishing the transmitters.
@@
Training on all samples. In the following, we consider the case of adversary $\mathcal{A}_{2}$ owning a model trained on both noise-free and noisy samples. Indeed, the adversary has been exposed to the full dataset (noise-free and noisy samples) associated with each device considered in our test. As before, we considered the Gaussian noise with standard deviation $\sigma=\{0,0.01,0.02,0.03,0.04,0.05\}$. Figure 10 shows the results of our analysis, where we perform a targeted attack (adversary $\mathcal{A}_{2}$) on each device of the pool (transmitter IDs). We report, for each of them, the MSE associated with the test set from the selected transmitter ID (green shaded area) and the MSE associated with the other measurements (other transmitters with any noise level). The probability distribution functions overlap for all the transmitter IDs. When considering the threshold defined by Eq. 6 (blue lines in Fig. 10), we obtain an average false negative ratio of about 0.66, confirming that the autoencoder fails to detect the presence of a specific transmitter in the radio spectrum when the training is performed on all the available samples (noise-free and noisy). Such a finding is confirmed by the empirical intuition that a model trained on all the samples cannot differentiate a specific transmitter from the pool since the (injected) noise makes the fingerprint of all the transmitters very similar to each other.
@@
Figure 10. MSE values when training on any samples (green curves) and testing on the rest of the dataset (red curves). The blue lines are the thresholds computed via Eq. 6. The overlap between the two sets of distributions (red and blue) is not negligible for all the transmitters (0.66 on average), and thus, the autoencoder fails to distinguish the transmitters.
7.3. Wireless Link
Our wireless scenario involves the same configuration as the wired one (recall Fig. 3). In line with the literature (alhazbi2023_acsac), we placed the transmitters and the receiver a few meters away in an office during working time, with people moving around in close proximity generating multipath. We consider the same data collection procedure as that described for Fig. 5, i.e., Gaussian noise type, CNN ResNet-18, and $10^{5}$ samples per image.
@@ 7.3. Wireless Link
Adversary $\mathcal{A}_{1}$. Figure 11 (top) shows the accuracy of ResNet-18 as a function of the noise standard deviation $\sigma$. When $\sigma=0$ (noise-free), the accuracy is approximately 0.96, while dropping to 0.16 when $\sigma=0.02$. It is worth noting that an accuracy value of 0.1 is equivalent to a random guess. The SNR is consistent with the one experienced when using the cable: we kept the attenuation to 30dB, and the distance between the transmitter and the receiver was only a few meters. Moreover, we observe the existence of a transient region in the range ]0,0.02]]0,0.02] where the accuracy is affected by the combination of different effects, i.e., the fingerprint of the device and the fingerprint of the noise introduced by the channel (sadighian2024ccnc; oligeri2024sac). The measurements involving the RF cable are not affected by external factors (multipath), and therefore, the fingerprint of the device is the principal component of the received signal. In contrast, when transmitting over the wireless channel, the channel impulse response varies over time depending on the communication link. Therefore, the DL model is likely to experience either consistent or different channel states (channel fingerprint). The final result, as observed by other contributions (sadighian2024ccnc; oligeri2024sac; alhazbi2023_acsac), is that RFF is affected by both the device and the channel fingerprint. The accuracy for $\sigma=0.01$ in Fig. 11 (top), i.e., 0.56, is higher than the one in Fig. 5 (top), i.e., 0.25, and it can be explained considering the higher impact of the wireless channel. Indeed, the measurements taken considering different transmitters have the same cable with the same channel fingerprint, while this is not happening in the wireless scenario, where different transmitters might be affected by a (slightly) different radio channel, which helps the classification. Finally, we observe that a minimum noise standard deviation of 0.02 fully anonymizes the transmitter in any scenario. Moreover, the SNR drop (0.1dB) necessary to remove the fingerprint is consistent with the one experienced using the cable.
@@
Figure 11. Adversary $\mathcal{A}_{1}$ training on noise-free samples from a wireless link. The adversary trains a model with 10 classes (one for each transmitter) on measurements with noise-free samples and then has to identify the transmitter(s) by considering measurements with noise strictly greater than zero.
@@
Adversary $\mathcal{A}_{2}$. We now consider adversary $\mathcal{A}_{2}$, i.e., a targeted attack against a specific transmitter in our pool of 10. We trained an autoencoder on wireless measurements taken from a specific device while setting the noise level $\sigma$ and tested it on the remainder of the measurements collected from the other transmitters in the pool, given the same $\sigma$. In fact, we assume $\mathcal{A}_{2}$ is capable of training from any type of measurement (noise-free and noisy) and then test it in the wild by collecting samples from the radio spectrum. Figure 12 shows the results of our analysis, where we consider the average Receiver Operating Characteristic (ROC) curves calculated considering the mean value over the 10 transmitters when varying $\sigma\in\{0,0.01,0.02,0.03,0.04,0.05\}$. Our results confirm previous findings. First, we observe that the performance of the autoencoder in a noise-free environment (solid blue line in Fig. 12)) is characterized by an FPR and TPR of 0.21 and 0.78, respectively, when considering the point in the ROC curve that offers the best trade-off between sensitivity and specificity, under the assumption that both types of classification errors (TPR and FPR) cost the same—the overall accuracy is equal to 0.81. Moreover, a Gaussian noise with standard deviation $\sigma=0.01$ is not enough to de-anonymize the transmitter. Indeed, the solid red line in Fig. 12, representing the ROC curve for $\sigma=0.01$, is characterized by FPR and TPR equal to about 0.27 and 0.7, respectively—the overall accuracy is equal to 0.73. Similarly to the previous case ($\mathcal{A}_{1}$), $\sigma=0.02$ allows us to achieve a much better anonymization level, i.e., TPR and FPR values equal to 0.27 and 0.7, respectively, with an overall accuracy equal to 0.64.
@@
Figure 12. Adversary $\mathcal{A}_{2}$. We trained an autoencoder on measurements taken from a target device and tested against the others in the pool while keeping the same noise level $\sigma\in\{0,0.01,0.02,0.03,0.04,0.05\}$ in an actual wireless link.
8. Comparison